Cost of debt using both methods Currently, Warren Industries can sell 15-year,
$1,000-par-value bonds paying annual interest at a 12% coupon rate. As a result of
current interest rates, the bonds can be sold for $1,010 each; flotation costs of $30
per bond will be incurred in this process. The firm is in the 40% tax bracket.
a. Find the net proceeds from sale of the bond, Nd.
b. Show the cash flows from the firm’s point of view over the maturity of the bond.
c. Calculate the before-tax and after-tax costs of debt.
d. Use the approximation formula to estimate the before-tax and after-tax costs of
debt.
e. Compare and contrast the costs of debt calculated in parts c and d. Which
approach do you prefer? Why?
1
Expert's answer
2021-07-20T15:30:57-0400

(a.) Net Proceeds = Sale price of Bond - Flotation cost

"= 1010 - 30\\

= 980"

(b)

At period "0= 1010-30=980"

between Period 1 to 15 = outflow of "1000\times0.12= - 120"

At the end of 15 years =1000

ie. Period        0      1-15        15

   CF            980    -120     -1000


c)

Before tax cost of Debt

"=RATE(15,120,-980,1000)= 12.30\%"

After Tax cost of Debt "=12.3 \times(1-0.4)=7.379\%=7.38\%"


d)

approximate method "= 120 + \frac{(1000-980)}{15} = 12.26\%"

 "=\frac{ (980+1000)}{2}"

after tax cost of debt "=12.26 \times(1-0.4)=7.36\%"


e)

The calculation method is more accurate than approximate method, however the results of approximate are fairly accurate but the calculation method is better.